OpenStudy (anonymous):

Write an equation of a line parallel to y = 2/3 x -5 in slope-intercept form that passes through the point (-6, -1).

OpenStudy (anonymous):

a. y= 2/3 x + 3 b. y= 2/3 x - 6 c. y= -3x - 1 d. y= 3x + 6

OpenStudy (anonymous):

i think it's B

OpenStudy (anonymous):

let us re-write the parallel equation using, y=mx+c I hope you are familiar with this

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

that too, for a parallel line, the m (slope/gradient) is the same so far our new equation is y=(2/3)x + b

OpenStudy (anonymous):

they said, x=-6 and y=-1 therefore we get -1=(2/3)(-1) + b we find for b,

OpenStudy (anonymous):

b the equation said was -5

OpenStudy (anonymous):

and x is -6

OpenStudy (anonymous):

@mickey4691